[EOD] no change in the general count highlighted in the [255pm] update.
[255pm] Count update -
Could be rushing the green count a bit, but here's the general profile.
[1110am] Some possibilities -
[815am] Overnight update -
On the ES, the black and green counts are now aligned, maximum potential is 1257.25/.75 based on these counts and assuming no extended wave (iii) / [iii].
In addition, the counts on the ES and YM are now identical.